An oscillator has an equation x(t) = (0.45 m) cos ( 359 t)

Assuming ω is in SI units, calculate the frequency in Hz to 2 sf.
Answer #1

x(t) = (0.45 m) cos ( 359 t)

compare it with  

x(t) =A*cos(ω*t)

so  ω = 359 rad/s

ω = 2*pi*f

f = ω /2pi = 359/(2*pi) =57.1366246 Hz

the frequency = 57 Hz answer

****************************************************************************
Goodluck for exam Comment in case any doubt, will reply for sure..